OpenStudy (anonymous):

athcalculus HELP: use your knowledge of the derivative to compute the limit given below: lim ((1/(x+h)^6)-(1/(x^6))/h h->0 the derivative that is being calculated is dy/dx

OpenStudy (anonymous):

is it -6/(x-h)?

OpenStudy (anonymous):

thinkg the quotient rule first?

OpenStudy (anonymous):

since it's a fraction

OpenStudy (anonymous):

have a look at this... you obviously have missed a few things in your class http://tutorial.math.lamar.edu/Classes/CalcI/DefnOfDerivative.aspx

OpenStudy (anonymous):

yes, i did.. it;s a summer session so we skipped a few.

OpenStudy (anonymous):

ps, i'm not bagging on you but am attempting to help. if you realize they've given you the definition of the derivative for the function f(x) = 1/x^6, then it's very easy to rewrite as f(x) = x^(-6) and then take the derivative

OpenStudy (anonymous):

the h is only in there because you using the definition and should not show up anywhere in your answer.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are being asked for the derivative of \(\frac{1}{x^6}\)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

?